Assume that y varies inversely as x. If y=8 when x=2.5, find x when y=2.
Answer by stanbon(75887)

Assume that y varies inversely as x. If y=8 when x=2.5, find x when y=2.
-----------------------
y = k/x
---
Find "k":
8 = k/2.5
k = 20
---
Equation:
y = 20/x
---
find x when y=2.
2 = 20/x
x = 10
================
Cheers,
Stan H.